A rectangle with a length of 8 has the same area as a trapezoid. One base of the trapezoid is 3 and the other base is three times the height of the rectangle. The height of the trapezoid is two thirds of the height of the rectangle. find the area of each of the two figures.
Answer by algebra hello(55)

let height of rectangle= h
then. 8*h=(1/2)(2 h/3)(3 3 h)
8= 1 h
h=7
area of rectangle= 8*7=56